### v1

@@ Abstract
Are we alone? It is a compelling question that human beings have confronted for centuries. The search for extraterrestrial life is a broad range of quests for finding the simple forms of life up to intelligent beings in the Universe. The plausible assumption is that there is a chance that intelligent life will followed by advanced civilization equipped or even dominated by artificial intelligence (AI). In this work, we categorize the advanced civilizations (on an equal footing, an AI-dominated civilization) on the Kardashev scale. We propose a new scale known as space exploration distance to measure civilization advancement. We propose a relation between this length and the Kardashev scale. Then, we suggest the idea that advanced civilizations will use primordial black holes as sources of harvesting energy. We calculate the energy harvested by calculating the space exploration distance. Finally, we propose an observational method to detect the possibility of extraterrestrial AI using Dyson spheres-like structures around primordial black holes in the Milky Way and other galaxies.
@@ 1 Introduction
Are we alone? It is a compelling question that human beings have confronted for centuries from antiquity to Giordano Bruno to modern times with the quest for finding extraterrestrial life by ”search for extraterrestrial intelligence” (SETI) project222see https://www.seti.org/. Nowadays, this question has become part of scientific investigation in a discipline known as astrobiology. Searching for life in the Universe is a multidisciplinary inquiry rooted in biology and astronomy (Plaxco & Gross, 2001). The expedition to find life can range from simple biological living systems to extraterrestrial intelligent life. Accordingly, searching for life is a vast ground of research starting from finding extra-solar planets (hereafter exoplanets) in habitable zones (Lee, 2018) to finding radio signals from intelligent life (Tarter, 2001; Rahvar, 2016). It is almost three decades from the breakthrough discovery of the first exoplanet (Mayor & Queloz, 1995). Nowadays, we find more than 5600 exoplanets  [^3] For the number of discovered and confirmed exoplanets, see https://science.nasa.gov/exoplanets/. The search for any bio-signature from the exoplanets is an extensive field of research. For this type of studies see Segura et al. (2005); Seager & Deming (2010); Schwieterman et al. (2018) and references therein.
@@
Parallel to these studies, the question of how extraterrestrial life could be intellectually complicated and intelligent becomes a burning question. In this direction, the Drake equation is a step forward to quantify the probability of finding intelligent life in the Universe (Drake, 1961). Three years after Drake’s insightful idea, Nikolai Kardashev, a soviet astronomer, attended an astronomy conference at Byurakan Observatory, Armenia, in 1964. At this meeting, Kardashev categorizes the three levels of civilization based on their consumption of energy (Kardashev, 1964). The first category is those who consume all the energy available to the planet. The second category uses all the energy of the host star by a method like the Dyson sphere  [^4] The Dyson sphere is a hypothetical space-based artificial megastructure around a star or other celestial energy source to collect effectively the energy for an advanced civilization. The first modern imagination of this structure is from a science fiction novel, ”Star Maker” by Olaf Stapledon in 1937 Dyson (1979). It is realized by a thought experiment presented by Freeman Dyson in his 1960 paper ”Search for Artificial Stellar Sources of Infrared Radiation”(Dyson, 1960a, b). (Dyson, 1960a, b), and the third level of civilization is related to the ones that use all the energy of the host galaxy.
@@
In this sequence, we are confronted with more advanced civilizations. In our case, on earth, as an example of a planet which hosts an intelligent life, it seems that after a century of the invention of Guglielmo Marconi, our species is on the verge of using artificial intelligence (AI) and the rise of AI is on the horizon. Due to the recent developments, there is a nonzero chance that AI will overtake and dominate the planet (Tegmark, 2017). In other words, the advanced civilizations become defined on the same footing as civilizations based on or dominated by AI  [^5] The result of this study is also validated if you replace the AI with an advanced civilization. Accordingly, the next question will be: What is the probability of finding ExtraTerrestrial - Artificial Intelligence (ET-AI) in the Universe? And how we can categorize them using the Kardashev scale. And what is our plan for the Search of the ExtraTerrestrial - Artificial Intelligence (SET-AI)?
@@
In this work, we introduce a new length scale as the space exploration distance, which is in a tight relation with the Kardashev scale. We assert that ET-AI will be on the Kardeshev scale between two and three. Accordingly, they explore space further from their host planet and star. Then, we estimate the energy harvested from the sources besides the host star. These extra sources could be the stellar or/and primordial black holes (PBH). The PBH is one of the compelling candidates for dark matter (DM) (Green & Kavanagh, 2021) and the most abundant and reliable energy source.
The structure of the work is as follows: In section 2, we introduce the theoretical background of this work. In section 3, we present our results on the search of ET-AI, and in the final section 4, we conclude.
@@ 2 Theoretical background
In this section, we review the Drake equation, the idea of using the Dyson sphere for harnessing energy, and the primordial black hole (PBH) as a plausible candidate of DM in three upcoming subsections.
2.1 Modified Drake equation for ET-AI
@@
where $R_{*}$ is the rate of star formation in solar masses per year, $f_{gs}$ is the fraction of the good stars, $f_{p}$ is the fraction of the stars that form planets, $n_{e}$ is the number of Earth-like planets, $f_{l}$ is the fraction of the planets where life originates, $f_{i}$ is the fraction planets that harbour intelligent life, $f_{c}$ is the fraction of planets that have intelligent life which has and desires to have interstellar communication and finally $L_{c}$ is the lifetime of the communicative intelligent species.
Now, for finding an ET-AI, even beyond our galaxy, using the idea of space mega-structures like Dyson spheres, we modify the Drake equation as below:
@@
where $N_{g}$ is the number of galaxies, which we define shortly. $f_{gg}$ is the fraction of good galaxies that can harbour life. Note that the star formation rate is a function of stellar mass and, in consequence, a function of galaxies’ total mass (Duarte Puertas et al., 2022). $f_{\rm{AI}}$ is the fraction of the planets that host intelligent life, and their dominant descendants are AI. $L_{\rm{AI}}$ is the lifespan of a typical AI-dominated civilization. As one may notice, we omit the probability of communicative intelligence fraction from the modified Drake equation 2. This is done because our search method for ET-AI is not based on radio communication, and probably the ET-AI are not social! We should note that two unknowns of $f_{gs}$ and $f_{gg}$ representing the ”good stars” and ”good galaxies” are themselves the topic of ongoing research (see Dayal et al. (2015); Whitmire (2020); Lacki (2021)). It is worth mentioning that most of the galaxies, due to their number of stars, are likely to harbour life, and we can set $f_{gg}\simeq 1$. However, the work by Dayal et al. (2015) is an example of a study to answer the question ”Which type of galaxy is most likely to host complex life in the cosmos?”. Coming back to the number of galaxies $N_{g}$, one knows that the preparation of the data and using the volume-limited constraints is a complicated survey-dependent procedure which reduces the number of observed galaxies in a specific redshift span (Yasuda et al., 2001). To obtain the number of galaxies $N_{g}$, which is calculated in the mass range of $(M_{i},M_{f})$ and redshift bin of $(z_{i},z_{f})$ we define:
@@
where $H(z)$ is the Hubble parameter, $D_{A}$ is the comoving angular diameter distance and $\Delta\Omega$ is the solid angle of a specific survey. $\frac{dn}{d\ln M}(M,z)$ is the comoving number density of dark matter halos (Press & Schechter, 1974; Cooray & Sheth, 2002) which host a central galaxy with the characteristics suitable for Drake equation. We should also note that at higher redshifts, the probability of finding AI/ advanced civilizations should be smaller due to the lack of time for the development of civilization. Additionally, as mentioned, the probability decreases because galaxies appear fainter and smaller at higher redshifts (both intrinsically and observationally). In conclusion, the modified Drake equation set the arena of searching ET-AI beyond our Milky Way.
@@ 2.2 Dyson spheres
The AI-based civilization, as a continuation of intelligent life, will be distinguished from its ancestors by the capability of using energy effectively and vastly. We assume that the ET-AI, at least, is in a second-level civilization due to the Kardashev scale. ET-AI uses space megastructures such as Dyson spheres for harnessing the energy of the host star (Dyson, 1960a, b) and beyond. On the other hand, an AI-based society is computational with its processing systems, which need a low temperature for high performance and, consequently, an enormous amount of energy. Accordingly, they will make space constructions like Dyson spheres at a distance from the host star or black hole to use all energy of the source at the desired temperature. These structures are heated up and will radiate in infrared (Wright et al., 2014). This idea motivates different groups to study the excess infrared radiation in the spectrum of stars for potential Dyson spheres (Contardo & Hogg, 2024; Suazo et al., 2024). On the other hand, they will use the shielding structures to bring the computational side of the structure to the temperature near interstellar medium in a typical galaxy.
The luminosity of a typical host star like the sun in the main sequence with a radius of $R_{\odot}\sim 7\times 10^{8}$ m is about $L_{\odot}\sim 3.8\times 10^{26}$ W. It is a good approximation that the Dyson sphere acts as a black body. Accordingly, the luminosity per unit area $j$ is related to temperature via Stephan-Boltzmann law:
@@
where $T$ is the temperature of the Black body in Kelvins and Stefan–Boltzmann constant is $\sigma\simeq 5.7\times 10^{-8}~{}{\text{W}/m^{2}{}\text{K}}^{4}$. So, the total power of the sun can be related to its temperature ($T_{\odot}\sim 5770$ K)
@@
More advanced civilizations start to explore the space beyond their star neighbourhood and harness the energy of the nearby stars. In this case, the typical radius of the Dyson sphere around a typical star $R_{\rm{Dyson}}|_{\rm{star}}$ can be approximated
@@
where $T_{\rm{Dyson}}$ is the black-body temperature of Dyson sphere. If we set the temperature of the Dyson sphere to $\sim 3000{\rm{K}}$ to avoid realistic solid material from melting, Dyson sphere-like space megastructures will be in the scale of $R_{\rm{Dyson}}|_{\rm{star}}\simeq 1.7\times 10^{-2}\rm{AU}$. For a temperature of $300{\rm{K}}$ for having liquid water or a classical computing regime the scale is $R_{\rm{Dyson}}|_{\rm{star}}\simeq 1.7\rm{AU}$ and for low-temperature computation around $30{\rm{K}}$ the scale will $R_{\rm{Dyson}}|_{\rm{star}}\simeq 170\rm{AU}$.
This range of temperatures suggests that we can find Dyson sphere-like structures in excess radiation in peak wavelength obtained from Wien’s law $\lambda_{\rm{max}}=\frac{b}{T}$, where $T$ is the absolute temperature and $b\simeq 2.898\times 10^{-3}{\rm{m.K}}$ is a constant of proportionality called Wien’s displacement constant.
Hsiao et al. (2021) shows that black holes (BH) can be more effective and promising energy sources than the main sequence stars. The major and main contribution from BH can be obtained from accretion disks. The matter in the accretion disk is heated up by friction due to the strong gravity of BH. For a black hole with mass $M$, the Eddington limit is below (Hsiao et al., 2021).
@@
where $m_{p}$ is the mass of proton, $\sigma_{T}=6.65\times 10^{-29}m^{-2}$ is the Thomson scattering cross section. The luminosity of the disk with the efficiency of $\eta_{\rm{disk}}$ and the accretion rate of $dm/dt$ is
@@
where for extreme Kerr BH we have $\eta_{\rm{disk}}=0.399$ and for non-rotating black holes $\eta_{\rm{disk}}=0.057$ (Thorne, 1974). The $L_{\rm{disk}}$ depends on the accretion rate, which can be in a range of $10^{-15}$ to $10^{-10}$ solar masses per second. For simplicity of calculations, we assume a BH (primordial) with a mass range of $0.01-100M_{\odot}$ with a conservative assumption of $\eta_{\rm{disk}}\sim 10^{-4}$, we set $L_{\rm{disk}}=10^{-4}L_{\rm{Edd}}$ which introduce power in order of $0.032-320L_{\odot}$. Black holes are categorized into families of stellar black holes, intermediate-mass black holes, supermassive black holes and primordial black holes. In the next subsection, we will briefly introduce PBHs as candidates for DM. As we did not observe the PBHs yet, one would question the assumptions, especially regarding the luminosity of their accretion disks and their mass, redshift and environment dependence. However, the assumption of an accretion disk around PBHs is discussed in various works see Ricotti, Ostriker, & Mack (2008); Ali-Haïmoud & Kamionkowski (2017); Safarzadeh (2018); Bhattacharjee & Stark (2020) and reference therein. As mentioned in our proposal, we choose the conservative accretion efficiency. Also, we assume that this efficiency is mass, redshift and environment independent.
Around a primordial black hole, one can make Dyson sphere-like structures closer due to the small size of the accretion disk, approximately in the distance of
@@
3000 Kelvin $\sim 0.017$ AU $\sim 0.03$ AU $\sim 966~{}{nm}$
300 Kelvin $\sim 1.7$ AU $\sim 3.0$ AU $\sim 9.66~{}\mu m$
30 Kelvin $\sim 170$ AU $\sim 300$ AU $\sim 96.6~{}\mu m$
Table 1: A table showing the Dyson sphere-like structures’ temperature (first column) and their distances from the energy source (star and PBH) in the second and third columns. The fourth column shows the peak of the wavelength in black-body radiation from a Dyson sphere-like structure. $R_{\rm{Dyson}}|_{\rm{\tiny{PBH}}}$ is calculated for solar mass PBH with $\eta_{\rm{disk}}=10^{-4}$.
In table 1, we list the distance of the Dyson sphere-like structure from the star or PBH corresponding to the desired temperature. In the final subsection, we discussed PBH.

### v2

@@ Abstract
Are we alone? It is a compelling question that human beings have confronted for centuries. The search for extraterrestrial life is a broad range of quests for finding simple forms of life up to intelligent beings in the Universe. The plausible assumption is that there is a chance that intelligent life will be followed by advanced civilization equipped or even dominated by artificial intelligence (AI). In this work, we categorize advanced civilizations (on an equal footing, an AI-dominated civilization) on the Kardashev scale. We propose a new scale known as the space exploration distance to measure civilization advancement. We propose a relation between this length and the Kardashev scale. Then, we suggest the idea that advanced civilizations will use primordial black holes (PBHs) as sources of harvesting energy. We calculate the energy harvested by calculating the space exploration distance. Finally, we propose an observational method to detect the possibility of extraterrestrial AI using Dyson sphere–like structures around PBHs in the Milky Way and other galaxies.
@@ 1 Introduction
Are we alone? It is a compelling question that human beings have confronted for centuries from antiquity to Giordano Bruno to modern times with the quest for finding extraterrestrial life by the ”search for extraterrestrial intelligence” project [^2] See https://www.seti.org/. Nowadays, this question has become part of scientific investigation in a discipline known as astrobiology. Searching for life in the Universe is a multidisciplinary inquiry rooted in biology and astronomy (Plaxco & Gross, 2001). The expedition to find life can range from simple biological living systems to extraterrestrial intelligent life. Accordingly, searching for life is a vast ground of research starting from finding extrasolar planets (hereafter exoplanets) in habitable zones (Lee, 2018) to finding radio signals from intelligent life (Tarter, 2001; Rahvar, 2016). It is almost three decades from the breakthrough discovery of the first exoplanet (Mayor & Queloz, 1995). Nowadays, we have found more than 5600 exoplanets  [^3] For the number of discovered and confirmed exoplanets, see https://science.nasa.gov/exoplanets/. The search for any biosignature from exoplanets is an extensive field of research. For this type of study see Segura et al. (2005); Seager & Deming (2010); Schwieterman et al. (2018) and references therein.
@@
Parallel to these studies, the question of how extraterrestrial life could be intellectually complicated and intelligent becomes a burning question. In this direction, the Drake equation is a step forward to quantify the probability of finding intelligent life in the Universe (Drake, 1961). Three years after Drake’s insightful idea, Nikolai Kardashev, a soviet astronomer, attended an astronomy conference at Byurakan Observatory, Armenia, in 1964. At this meeting, Kardashev categorized the three levels of civilization based on their consumption of energy (Kardashev, 1964). The first category is those who consume all the energy available to the planet. The second category uses all the energy of the host star by a method like a Dyson sphere  [^4] The Dyson sphere is a hypothetical space-based artificial megastructure around a star or other celestial energy source, which is used to collect effectively the energy for an advanced civilization. The first modern imagination of this structure is from a science fiction novel, ”Star Maker” by Olaf Stapledon in 1937 Dyson (1979). It is realized by a thought experiment presented by Freeman Dyson in his 1960 paper ”Search for Artificial Stellar Sources of Infrared Radiation”(Dyson, 1960a, b). (Dyson, 1960a, b), and the third level of civilization is related to the ones that use all the energy of the host galaxy.
@@
In this sequence, we are confronted with more advanced civilizations. In our case, on Earth, as an example of a planet which hosts intelligent life, it seems that after a century of the invention of Guglielmo Marconi, our species is on the verge of using artificial intelligence (AI) and the rise of AI is on the horizon. Due to recent developments, there is a nonzero chance that AI will overtake and dominate the planet (Tegmark, 2017). In other words, advanced civilizations become defined on the same footing as civilizations based on or dominated by AI  [^5] The result of this study is also validated if you replace the AI with an advanced civilization. Accordingly, the next question will be: what is the probability of finding extraterrestrial artificial intelligence (ET-AI) in the Universe? How we can categorize it using the Kardashev scale. And what is our plan for the search for extraterrestrial artificial intelligence (SET-AI)?
@@
In this work, we introduce a new length scale as the space exploration distance (SED), which is in a tight relation with the Kardashev scale. We assert that ET-AI will be on the Kardeshev scale between two and three. Accordingly, explore is explored further from the host planet and star. Then, we estimate the energy harvested from the sources besides the host star. These extra sources could be the stellar or/and primordial black holes (PBH). PBHs are one of the compelling candidates for dark matter (DM) (Green & Kavanagh, 2021) and the most abundant and reliable energy source.
The structure of the work is as follows: In section 2, we introduce the theoretical background of this work. In section 3, we present our results on SET-AI, and in the final section 4, we conclude.
@@ 2 Theoretical background
In this section, we review the Drake equation, the idea of using a Dyson sphere for harnessing energy, and PBHs as a plausible candidate of DM in three upcoming subsections.
2.1 Modified Drake equation for Extraterrestrial Artificial Intelligence
@@
where $R_{*}$ is the rate of star formation in solar masses per year, $f_{gs}$ is the fraction of the good stars, $f_{p}$ is the fraction of the stars that form planets, $n_{e}$ is the number of Earth-like planets, $f_{l}$ is the fraction of the planets where life originates, $f_{i}$ is the fraction planets that harbor intelligent life, $f_{c}$ is the fraction of planets that have intelligent life which has and desires to have interstellar communication, and finally $L_{c}$ is the lifetime of the communicative intelligent species.
Now, for finding an ET-AI, even beyond our Galaxy, using the idea of space megastructures like Dyson spheres, we modify the Drake equation as below
@@
where $N_{g}$ is the number of galaxies, which we define shortly. $f_{gg}$ is the fraction of good galaxies that can harbor life. Note that the star formation rate is a function of stellar mass and, in consequence, a function of the galaxy’s total mass (Duarte Puertas et al., 2022). $f_{\rm{AI}}$ is the fraction of planets that host intelligent life, and their dominant descendants are AI. $L_{\rm{AI}}$ is the lifespan of a typical AI-dominated civilization. As one may notice, we omit the probability of communicative intelligence fraction from the modified Drake equation 2. This is done because our search method for ET-AI is not based on radio communication, and probably ET-AI are not social! We should note that two unknowns of $f_{gs}$ and $f_{gg}$ representing ”good stars” and ”good galaxies” are themselves a topic of ongoing research (see Dayal et al. (2015); Whitmire (2020); Lacki (2021)). It is worth mentioning that most of the galaxies, due to their numbers of stars, are likely to harbor life, and we can set $f_{gg}\simeq 1$. However, the work by Dayal et al. (2015) is an example of a study to answer the question ”which type of galaxy is most likely to host complex life in the cosmos?” Coming back to the number of galaxies $N_{g}$, one knows that the preparation of the data and using volume-limited constraints is a complicated survey-dependent procedure which reduces the number of observed galaxies in a specific redshift span (Yasuda et al., 2001). To obtain the number of galaxies $N_{g}$, which is calculated in the mass range of $(M_{i},M_{f})$ and redshift bin of $(z_{i},z_{f})$ we define
@@
where $H(z)$ is the Hubble parameter, $D_{A}$ is the comoving angular diameter distance, and $\Delta\Omega$ is the solid angle of a specific survey. $\frac{dn}{d\ln M}(M,z)$ is the comoving number density of DM halos (Press & Schechter, 1974; Cooray & Sheth, 2002) which host a central galaxy with the characteristics suitable for Drake equation. We should also note that at higher redshifts, the probability of finding AI/advanced civilizations should be smaller due to the lack of time for the development of civilization. Additionally, as mentioned, the probability decreases because galaxies appear fainter and smaller at higher redshifts (both intrinsically and observationally). In conclusion, the modified Drake equation sets the arena of SET-AI beyond our Milky Way.
@@ 2.2 Dyson spheres
An AI-based civilization, as a continuation of intelligent life, will be distinguished from its ancestors by the capability of using energy effectively and vastly. We assume that the ET-AI, at least, is in a second-level civilization on the Kardashev scale. ET-AI uses space megastructures such as Dyson spheres for harnessing the energy of their host star (Dyson, 1960a, b) and beyond. On the other hand, an AI-based society is computational with processing systems, which need a low temperature for high performance and, consequently, an enormous amount of energy. Accordingly, they will make space constructions like Dyson spheres at a distance from the host star or black hole to use all energy of the source at the desired temperature. These structures are heated up and will radiate in infrared (Wright et al., 2014). This idea motivates different groups to study excess infrared radiation in the spectrum of stars as potential Dyson spheres (Contardo & Hogg, 2024; Suazo et al., 2024). On the other hand, they will use the shielding structures to bring the computational side of the structure to the temperature near interstellar medium in a typical galaxy.
The luminosity of a typical host star like the sun on the main sequence with a radius of $R_{\odot}\sim 7\times 10^{8}$ m is about $L_{\odot}\sim 3.8\times 10^{26}$ W. It is a good approximation that a Dyson sphere acts as a blackbody. Accordingly, the luminosity per unit area $j$ is related to temperature via Stephan-Boltzmann law:
@@
where $T$ is the temperature of the blackbody in kelvins and the Stefan–Boltzmann constant is $\sigma\simeq 5.7\times 10^{-8}~{}{\text{W}/m^{2}{}\text{K}}^{4}$. So, the total power of the sun can be related to its temperature ($T_{\odot}\sim 5770$ K) as
@@
More advanced civilizations start to explore the space beyond their stellar neighborhood and harness the energy of nearby stars. In this case, the typical radius of the Dyson sphere around a typical star $R_{\rm{Dyson}}|_{\rm{star}}$ can be approximated
@@
where $T_{\rm{Dyson}}$ is the blackbody temperature of the Dyson sphere. If we set the temperature of the Dyson sphere to $\sim 3000~{}{\rm{K}}$ to avoid realistic solid material from melting, Dyson sphere–like space megastructures will be on the scale of $R_{\rm{Dyson}}|_{\rm{star}}\simeq 1.7\times 10^{-2}\rm{AU}$. For a temperature of $300~{}{\rm{K}}$ for having liquid water or a classical computing regime the scale is $R_{\rm{Dyson}}|_{\rm{star}}\simeq 1.7\rm{AU}$ and for low-temperature computation around $30{\rm{K}}$ the scale will $R_{\rm{Dyson}}|_{\rm{star}}\simeq 170\rm{AU}$.
This range of temperatures suggests that we can find Dyson sphere–like structures as excess radiation in the peak wavelength obtained from Wien’s law $\lambda_{\rm{max}}=\frac{b}{T}$, where $T$ is the absolute temperature and $b\simeq 2.898\times 10^{-3}{\rm{mK}}$ is a constant of proportionality called Wien’s displacement constant.
Hsiao et al. (2021) shows that BHs can be a more effective and promising energy sources than the main-sequence stars. The major and main contribution from a BH can be obtained from accretion disks. The matter in the accretion disk is heated up by friction due to the strong gravity of the BH. For a BH with mass $M$, the Eddington limit is (Hsiao et al., 2021)
@@
where $m_{p}$ is the mass of the proton and $\sigma_{T}=6.65\times 10^{-29}~{}{\rm{m}}^{-2}$ is the Thomson scattering cross section. The luminosity of the disk with the efficiency of $\eta_{\rm{disk}}$ and the accretion rate of $dm/dt$ is
@@
where for an extreme Kerr BH we have $\eta_{\rm{disk}}=0.399$ and for a nonrotating BH $\eta_{\rm{disk}}=0.057$ (Thorne, 1974). $L_{\rm{disk}}$ depends on the accretion rate, which can be in the range of $10^{-15}-10^{-10}M_{\odot}s^{-1}$. For simplicity of calculation, we assume a PBH with a mass range of $0.01-100M_{\odot}$ with a conservative assumption of $\eta_{\rm{disk}}\sim 10^{-4}$, and we set $L_{\rm{disk}}=10^{-4}L_{\rm{Edd}}$ which introduce powers of the order of $0.032-320L_{\odot}$. BHs are categorized into families of stellar BHs, intermediate-mass BHs, supermassive BHs. In the next subsection, we will briefly introduce PBHs as a candidates for DM. As we have not observed PBHs yet, one would question the assumptions, especially regarding the luminosity of their accretion disks and their mass, redshift, and environment dependence. However, the assumption of an accretion disk around PBHs is discussed in various works, see Ricotti, Ostriker, & Mack (2008); Ali-Haïmoud & Kamionkowski (2017); Safarzadeh (2018); Bhattacharjee & Stark (2020) and references therein. As mentioned in our proposal, we choose a conservative accretion efficiency. Also, we assume that this efficiency is mass, redshift, and environment independent.
Around a PBH, one can make Dyson sphere–like structures closer due to the small size of the accretion disk, approximately in the distance of
@@
3000 K $\sim 0.017$ AU $\sim 0.03$ AU $\sim 966~{}{\rm{nm}}$
300 K $\sim 1.7$ AU $\sim 3.0$ AU $\sim 9.66~{}\mu{\rm{m}}$
30 K $\sim 170$ AU $\sim 300$ AU $\sim 96.6~{}\mu{\rm{m}}$
Table 1: A table showing the Dyson sphere–like structures’ temperature (first column) and their distances from the energy source (star and Primordial Black Hole) in the second and third columns. The fourth column shows the peak of the wavelength in blackbody radiation from a Dyson sphere–like structure. $R_{\rm{Dyson}}|_{\rm{\tiny{PBH}}}$ is calculated for a solar mass PBH with $\eta_{\rm{disk}}=10^{-4}$.
In Table 1, we list the distances of Dyson sphere–like structures from the star or PBH corresponding to the desired temperature. In the final subsection, we discussed PBHs.